OpenStudy (anonymous):

If the coordinates of the points D, E, G, and H are −3, −6, 9, and 11, respectively, which segment is longest? Segment DE Segment EH Segment GH Segment DH

OpenStudy (anonymous):

EH

OpenStudy (anonymous):

DH=6+11=17

OpenStudy (anonymous):

DE = -6-(-3) = 3 units EH = 9-(-6) = 15 units GH = 11-9 = 2 units DH = 11-(-3) = 13 units hence EH is the longest segment

OpenStudy (anonymous):

thank you(:

OpenStudy (anonymous):

EH=11-(16)=17